OpenStudy (anonymous):

find the zero of f(x)3/4x-12

OpenStudy (anonymous):

\[f(x)=\frac{ 3 }{ 4 }x+12\]

OpenStudy (anonymous):

its - 12 sorry

OpenStudy (anonymous):

i got x=16

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }x-12=0\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

thank you though :3